# On a conjecture of Peter Neumann on fixed points in permutation groups

### v1

@@ 9.3. The proof
Assume now $S$ is one of $\mathrm{PSU}_{7}(2),\mathrm{PSp}_{4}(7),\mathrm{PSp}_{8}(3),F_{4}(2),Co_{1},Ly,J_{4}$. The group $S=Co_{1}$ was handled above in the case $t\geq 4$ (and we did not use this assumption). In order to handle the other groups, we keep the notation $t$, $V_{0}$ and $U$ as above. It is sufficient to prove that $t\geq 4$, as this reduces to a case already considered; also we may assume $f>n/10$ and in particular $n<220$. If $S=\mathrm{PSU}_{7}(2)$ then by [32] we have $n=42,43$. We have $|g|=\Phi^{*}_{14}(2)=43$, and so $t\geq 4$ (since otherwise $\dim(V_{0})\leq 3f\leq 21$; but all Jordan blocks on a complement have size at least $43$). If $S=\mathrm{PSp}_{4}(7)$ then by [32] we have $n=24,25$ or $n\geq 126$. Since $f=4$ we may assume $n=24,25$. But $|g|=\Phi^{*}_{4}(7)=25$, so as in the previous case we have $t\geq 4$ and we are done. Assume now $S=\mathrm{PSp}_{8}(3)$. Then by [32] we have $n=40,41$. Since $|g|=\Phi^{*}_{8}(3)=41$ we deduce $t\geq 4$ exactly as above. Assume $S=F_{4}(2)$, so $n=52$ and $|g|=13$. We have $G=\Omega^{\varepsilon}_{n}(q)$. If $t\leq 3$ then $\dim(V_{0})\leq 3f=36$. But all Jordan blocks on a complement have size $13$, so in fact there are at least two such blocks, and so $\dim(V_{0})\leq 26$. But also $a=\dim(C_{V}(g))\leq n/2=26$ and all Jordan blocks on $V_{0}$ have the same size, so $d\leq 2\cdot 15^{2}+11\cdot 2^{2}=494$ (attained if $g$ has two Jordan blocks of size $13$ and the others of size $2$), which implies $|C_{G}(g)|<|G:M|^{1/3}$. If $S=Ly$ or $J_{4}$, then since $p$ divides $|g|$, by [32] we see that there are no cases with $n<220$. The proof is finally concluded. ∎
@@
    $S=\mathrm{P}\Omega_{7}(q)$, $\mathrm{P}\Omega_{9}(q)$ or $\mathrm{P}\Omega^{+}_{10}(q)$ and $V$ is a spin module of dimension $8$, $16$ or $16$ (here we allow $q$ even for $\mathrm{P}\Omega_{7}(q)$ and $\mathrm{P}\Omega_{9}(q)$);
@@
(iii) For $n=8$ we have $G=\Omega^{+}_{8}(q)$, and $S$ is conjugate in $\text{Aut}(G/Z(G))$ to the stabilizer of a nonsingular vector, which has been handled already. For $n=16$ and $S=\mathrm{P}\Omega_{9}(q)$ or $\mathrm{P}\Omega^{+}_{10}(q)$, the restriction of $V$ to $\text{Spin}^{+}_{8}(q)$ is the sum of the two half-spin modules of dimension $8$, so by letting $g$ be an element of order $q^{4}-1$, we have that each eigenspace of $g$ has dimension at most $2$, and the result follows.
@@
Assume first $S=\mathrm{PSL}_{2}(r)$. We may assume $r\geq 7$ and $r\neq 9$ in view of $\mathrm{PSL}_{2}(4)\cong\mathrm{PSL}_{2}(5)\cong A_{5}$ and $\mathrm{PSL}_{2}(9)\cong A_{6}$. Suppose $p\mid r$, that is, $S\in\text{Lie}(p)$. If $V$ is $p$-restricted as $L$-module, then $r=q$, $V$ is a symmetric power of the natural 2-dimensional module, and a regular unipotent element $g$ of $L$ is regular in $\mathrm{SL}_{n}(q)$. Moreover, $g^{G}$ splits into at most (2,q1)(2,q-1)-classes, and by [7] we have $n\geq 4$; it follows that $\mathrm{fix}(g,G/M)\leq(2,q-1)|C_{G}(g)|/q<|G:M|^{1/3}$. For example, for $n=4$ we have $G=\mathrm{Sp}_{4}(q)$, $p\geq 5$, and $|C_{G}(g)|=(2,q-1)q^{2}$, from which $\mathrm{fix}(g,G/M)\leq(2,q-1)^{2}q<|G:M|^{1/3}$. If $V$ is not $p$-restricted, then by [70] the possibilities are $\mathrm{PSL}_{2}(q^{f})\cong\mathrm{PSp}_{2}(q^{f})<\mathrm{PSp}_{2^{f}}(q)$ with $qf$ odd, and $\mathrm{PSL}_{2}(q^{f})\cong\Omega_{3}(q^{f})<\Omega_{3^{f}}(q)$ with $qf$ odd. We have $V=W\otimes W^{(q)}\otimes\cdots\otimes W^{(q^{f-1})}$ where $\dim(W)=2$ or $\dim(W)=3$ in the respective cases. We let $g\in L$ be of order $\Phi^{*}_{2}(q^{f})$. Assume first $n=8$. We see that $g$ has an eigenvalue generating $\mathbb{F}_{q^{6}}$, so $g$ has six distinct nontrivial eigenvalues and the conclusion follows. In all the other cases, we see that $g$ has distinct eigenvalues on $W\otimes W^{(q)}$. In particular, each eigenspace on $V$ has dimension at most $\dim(W)^{f-2}$, which is sufficient.
@@
Assume now $S=\mathrm{PSL}^{\pm}_{3}(r)$ with $r\geq 3$. We choose $|g|=\Phi^{*}_{3}(r)$ for $\mathrm{PSL}_{3}(r)$ and $|g|=\Phi^{*}_{6}(r)$ for $\mathrm{PSU}_{3}(r)$, so $N:=N_{L}(\langle g\rangle)/C_{L}(g)$ acts with all orbits of size $3$ on the set of nontrivial eigenspaces of $g$ on $V$. Note also that by [27, Propositions 3.11 and 3.13] $S$ is generated by two conjugates of $g$, so $a=\dim(C_{V}(g))\leq n/2$.
@@
Suppose $x=y$, that is, $V$ is self-dual. Then, for each $N$-orbit $\mathcal{O}$ of nontrivial eigenspaces, there is another orbit whose eigenspaces have the same dimension (and whose corresponding eigenvalues are inverses of those in $\mathcal{O}$). Since $a=\dim(C_{V}(g))\leq n/2$, we may apply Lemmas 9.3 and 9.4 with $c=1/2$ and $B=6$ to deduce:
@@
If $n$ is at least these values, then by Lemmas 9.3 and 9.4 we have $|C_{G}(g)|<|G:M|^{1/3}$. Assume then $n$ is smaller than these values; by [61], the only remaining possibility is the adjoint module for $S$, with $n=7$ or $8$ according to whether $p=3$ or $p\neq 3$. For $n=7$ we have $M\not\in A$ by [7]. For $n=8$ we have $a=2$, and as above every nontrivial eigenspace of $g$ has dimension $1$. Moreover $G=\Omega^{\pm}_{8}(q)$; by Lemma 7.10, the dimension of the centralizer of $g$ in $\mathrm{SO}_{8}(K)$ is $4$ and we see that $|C_{G}(g)|<|G:M|^{1/3}$.
@@
Assume $V$ is $p$-restricted as $L$-module. It follows from Premet’s theorem [68, Theorem 1] that every dominant weight $\gamma=(c,d)$ with $\gamma\in X:=\lambda-\mathbb{Z}\alpha_{1}-\mathbb{Z}\alpha_{2}$ is a weight of $V$. Suppose $x\geq y+2$. Since $(x-y,0)=\lambda-(y,y)=\lambda-y\alpha_{1}-y\alpha_{2}\in X$, again by Premet’s theorem, $V$ contains all the weights corresponding to the representation with highest weight (xy,0)(x-y,0), which is the (xy)(x-y)-th symmetric power. Since $x-y\geq 2$, and since $g$ has at least six distinct nontrivial eigenvalues on $(2,0)$, we deduce that $g$ has at least six distinct nontrivial eigenvalues on $V$, as claimed. Now suppose $x=y+1$. Then $(2,1)=\lambda-(y-1,y-1)\in X$, so $V$ contains all the weights corresponding to the representation with highest weight $(2,1)$. This has dimension $15$ and is a constituent of $S^{2}(W)\otimes W^{*}$ where $W$ is the $3$-dimensional natural module. We can then calculate that $g$ has at least six distinct nontrivial eigenvalues on $V$.
@@
Denote by $E$ the number of distinct eigenvalues of $g$. Since $a\leq n/2$ and since nontrivial eigenspaces occur in orbits of size $3$, it can be seen that $d+E$ is at most
@@
(attained when $g$ has trivial eigenspace of dimension $n/2$, three eigenspaces of dimension $1$ and three eigenspaces of dimension (n/23)/3(n/2-3)/3), from which it also follows that $d\leq n^{2}/3-n+6$. If $G=\mathrm{SL}_{n}(q)$ then we deduce $|C_{G}(g)|^{3}<q^{n^{2}-3n+18}$, which is <|G:M|<|G:M| if $n\geq 15$. If $G=\mathrm{SU}_{n}(q^{1/2})$ then, setting $q_{0}=q^{1/2}$, by Lemma 7.11 we deduce $|C_{G}(g)|^{3}<q_{0}^{3(d+E)}\leq q_{0}^{n^{2}-3n+39}$, which is <|G:M|<|G:M| if $n\geq 24$.
Let us address the remaining cases (i.e. the case $n\leq 14$ for $G=\mathrm{SL}_{n}(q)$ and the case $n\leq 23$ for $G=\mathrm{SU}_{n}(q^{1/2})$). Assume first $V$ is $p$-restricted; the options are $n=6,10$ for $G=\mathrm{SL}_{n}(q)$; and $n=6,10,15,18,21$ for $G=\mathrm{SU}_{n}(q^{1/2})$. If $n=6$ then $V=S^{2}(W)$ (where $W$ is the natural 3-dimensional module), $g$ is regular and we have $|C_{G}(g)|<|G:M|^{1/3}$. If $n=10$ then $V=S^{3}(W)$; we see that $a=\dim(C_{V}(g))=1$ and so $|C_{G}(g)|<|G:M|^{1/3}$. If $n=15$ then it must be $a\leq 6$, which implies $d\leq 51$ and the result follows. If $n=18$ then $p=5$ and $V$ is a constituent of $S^{3}(W)\otimes W^{*}$. Since by the above the largest eigenspace on $S^{3}(W)$ has dimension at most $2$, it follows that $a\leq 6$, which gives the conclusion. Finally, if $n=21$ then $a\leq 9$, so $d\leq 111=3\cdot 6^{2}+3$ and this is sufficient. Assume now $V$ is not $p$-restricted with $n\leq 23$; by [70], the possible embeddings are of type $\mathrm{SL}_{3}(q^{2})<\mathrm{SL}_{9}(q)$ and $\mathrm{SL}_{3}(q)<\mathrm{SU}_{9}(q^{1/2})$. In the first case, $a=3$ and every nontrivial eigenspace has dimension $1$, so $d=15$ and the result follows. In the second case, there are three nontrivial eigenspaces of dimension $1$, and three nontrivial eigenspaces of dimension $2$, so again $d=15$ and $|C_{G}(g)|<|G:M|^{1/3}$ also in this case.
@@
Suppose $g$ is semisimple. Note that every prime divisor of $|g|$ is at least $7$ and also we have $|N_{L}(\langle g\rangle)/C_{L}(g)|=3$. By [80, Theorem 1.2], we deduce that $g$ has at least $5$ distinct eigenvalues. Since these are permuted in orbits of size $3$, we deduce that $g$ has at least $6$ distinct eigenvalues. In particular, by the same calculation as in the case $p\mid r$ we see that $d\leq n^{2}/3-n+6$ and $d+E\leq n^{2}/3-n+13$, where $E$ denotes the number of distinct eigenvalues of $g$. Denote also by $E_{2}$ the number of distinct irreducible factors of degree at least two of the characteristic polynomial of $g$. Next we upper bound $|C_{G}(g)|$ using Lemmas 7.10 and 7.11.
@@
Suppose finally $g$ is not semisimple; we apply Lemma 7.4. Letting $P$ be a Sylow $p$-subgroup of $\langle g\rangle$, we have that $V\downarrow P=V_{0}\oplus U$ where all Jordan blocks on $V_{0}$ have size $|P|\geq 7$, and $U$ is the sum of at most $3$ indecomposable module of dimension $t$. If $t\geq 4$ then applying Lemma 7.8 with $C=0$ and $B=4$ we deduce $d\leq n^{2}/4$. In particular, the exact same computation as in Lemma 9.8 reduces to the case $n\leq 29$, for which we are done (as we are assuming $n\geq 41$).

### v2

@@ 9.3. The proof
Assume now $S$ is one of $\mathrm{PSU}_{7}(2),\mathrm{PSp}_{4}(7),\mathrm{PSp}_{8}(3),F_{4}(2),Co_{1},Ly,J_{4}$. The group $S=Co_{1}$ was handled above in the case $t\geq 4$ (and we did not use this assumption). In order to handle the other groups, we keep the notation $t$, $V_{0}$ and $U$ as above. It is sufficient to prove that $t\geq 4$, as this reduces to a case already considered; also we may assume $f>n/10$ and in particular $n<220$. If $S=\mathrm{PSU}_{7}(2)$ then by [32] we have $n=42,43$. We have $|g|=\Phi^{*}_{14}(2)=43$, and so $t\geq 4$ (since otherwise $\dim(V_{0})\leq 3f\leq 21$; but all Jordan blocks on a complement have size at least $43$). If $S=\mathrm{PSp}_{4}(7)$ then by [32] we have $n=24,25$ or $n\geq 126$. Since $f=4$ we may assume $n=24,25$. But $|g|=\Phi^{*}_{4}(7)=25$, so as in the previous case we have $t\geq 4$ and we are done. Assume now $S=\mathrm{PSp}_{8}(3)$. Then by [32] we have $n=40,41$. Since $|g|=\Phi^{*}_{8}(3)=41$ we deduce $t\geq 4$ exactly as above. Assume $S=F_{4}(2)$, so $n=52$ and $|g|=13$. We have $G=\Omega^{\varepsilon}_{n}(q)$. If $t\leq 3$ then $\dim(V_{0})\leq 3f=36$. But all Jordan blocks on a complement have size $13$, so in fact there are at least two such blocks, and so $\dim(V_{0})\leq 26$. But also $a=\dim(C_{V}(g))\leq n/2=26$ and all Jordan blocks on $V_{0}$ have the same size, so $d\leq 2\cdot 15^{2}+11\cdot 2^{2}=494$ (attained if $g$ has two Jordan blocks of size $13$ and the others of size 22), which implies $|C_{G}(g)|<|G:M|^{1/3}$. If $S=Ly$ or $J_{4}$, then since $p$ divides $|g|$, by [32] we see that there are no cases with $n<220$. The proof is finally concluded. ∎
@@
    $S=\mathrm{P}\Omega_{7}(q)$, $\mathrm{P}\Omega_{9}(q)$ or $\mathrm{P}\Omega^{+}_{10}(q)$ and $V$ is a spin module of dimension 88, $16$ or $16$ (here we allow $q$ even for $\mathrm{P}\Omega_{7}(q)$ and $\mathrm{P}\Omega_{9}(q)$);
@@
(iii) For $n=8$ we have $G=\Omega^{+}_{8}(q)$, and $S$ is conjugate in $\text{Aut}(G/Z(G))$ to the stabilizer of a nonsingular vector, which has been handled already. For $n=16$ and $S=\mathrm{P}\Omega_{9}(q)$ or $\mathrm{P}\Omega^{+}_{10}(q)$, the restriction of $V$ to $\text{Spin}^{+}_{8}(q)$ is the sum of the two half-spin modules of dimension 88, so by letting $g$ be an element of order $q^{4}-1$, we have that each eigenspace of $g$ has dimension at most 22, and the result follows.
@@
Assume first $S=\mathrm{PSL}_{2}(r)$. We may assume $r\geq 7$ and $r\neq 9$ in view of $\mathrm{PSL}_{2}(4)\cong\mathrm{PSL}_{2}(5)\cong A_{5}$ and $\mathrm{PSL}_{2}(9)\cong A_{6}$. Suppose $p\mid r$, that is, $S\in\text{Lie}(p)$. If $V$ is $p$-restricted as $L$-module, then $r=q$, $V$ is a symmetric power of the natural 2-dimensional module, and a regular unipotent element $g$ of $L$ is regular in $\mathrm{SL}_{n}(q)$. Moreover, $g^{G}$ splits into at most (2,q1)(2,q-1)-classes, and by [7] we have $n\geq 4$; it follows that $\mathrm{fix}(g,G/M)\leq(2,q-1)|C_{G}(g)|/q<|G:M|^{1/3}$. For example, for $n=4$ we have $G=\mathrm{Sp}_{4}(q)$, $p\geq 5$, and $|C_{G}(g)|=(2,q-1)q^{2}$, from which $\mathrm{fix}(g,G/M)\leq(2,q-1)^{2}q<|G:M|^{1/3}$. If $V$ is not $p$-restricted, then by [70] the possibilities are $\mathrm{PSL}_{2}(q^{f})\cong\mathrm{PSp}_{2}(q^{f})<\mathrm{PSp}_{2^{f}}(q)$ with qfqf odd, and $\mathrm{PSL}_{2}(q^{f})\cong\Omega_{3}(q^{f})<\Omega_{3^{f}}(q)$ with qfqf odd. We have $V=W\otimes W^{(q)}\otimes\cdots\otimes W^{(q^{f-1})}$ where $\dim(W)=2$ or $\dim(W)=3$ in the respective cases. We let $g\in L$ be of order $\Phi^{*}_{2}(q^{f})$. Assume first $n=8$. We see that $g$ has an eigenvalue generating $\mathbb{F}_{q^{6}}$, so $g$ has six distinct nontrivial eigenvalues and the conclusion follows. In all the other cases, we see that $g$ has distinct eigenvalues on $W\otimes W^{(q)}$. In particular, each eigenspace on $V$ has dimension at most $\dim(W)^{f-2}$, which is sufficient.
@@
Assume now $S=\mathrm{PSL}^{\pm}_{3}(r)$ with $r\geq 3$. We choose $|g|=\Phi^{*}_{3}(r)$ for $\mathrm{PSL}_{3}(r)$ and $|g|=\Phi^{*}_{6}(r)$ for $\mathrm{PSU}_{3}(r)$, so $N:=N_{L}(\langle g\rangle)/C_{L}(g)$ acts with all orbits of size 33 on the set of nontrivial eigenspaces of $g$ on $V$. Note also that by [27, Propositions 3.11 and 3.13] $S$ is generated by two conjugates of $g$, so $a=\dim(C_{V}(g))\leq n/2$.
@@
Suppose $x=y$, that is, $V$ is self-dual. Then, for each NN-orbit $\mathcal{O}$ of nontrivial eigenspaces, there is another orbit whose eigenspaces have the same dimension (and whose corresponding eigenvalues are inverses of those in $\mathcal{O}$). Since $a=\dim(C_{V}(g))\leq n/2$, we may apply Lemmas 9.3 and 9.4 with $c=1/2$ and $B=6$ to deduce:
@@
If $n$ is at least these values, then by Lemmas 9.3 and 9.4 we have $|C_{G}(g)|<|G:M|^{1/3}$. Assume then $n$ is smaller than these values; by [61], the only remaining possibility is the adjoint module for $S$, with $n=7$ or 88 according to whether $p=3$ or $p\neq 3$. For $n=7$ we have $M\not\in A$ by [7]. For $n=8$ we have $a=2$, and as above every nontrivial eigenspace of $g$ has dimension 11. Moreover $G=\Omega^{\pm}_{8}(q)$; by Lemma 7.10, the dimension of the centralizer of $g$ in $\mathrm{SO}_{8}(K)$ is 44 and we see that $|C_{G}(g)|<|G:M|^{1/3}$.
@@
Assume $V$ is $p$-restricted as $L$-module. It follows from Premet’s theorem [68, Theorem 1] that every dominant weight $\gamma=(c,d)$ with $\gamma\in X:=\lambda-\mathbb{Z}\alpha_{1}-\mathbb{Z}\alpha_{2}$ is a weight of $V$. Suppose $x\geq y+2$. Since $(x-y,0)=\lambda-(y,y)=\lambda-y\alpha_{1}-y\alpha_{2}\in X$, again by Premet’s theorem, $V$ contains all the weights corresponding to the representation with highest weight (xy,0)(x-y,0), which is the (xy)(x-y)-th symmetric power. Since $x-y\geq 2$, and since $g$ has at least six distinct nontrivial eigenvalues on $(2,0)$, we deduce that $g$ has at least six distinct nontrivial eigenvalues on $V$, as claimed. Now suppose $x=y+1$. Then $(2,1)=\lambda-(y-1,y-1)\in X$, so $V$ contains all the weights corresponding to the representation with highest weight $(2,1)$. This has dimension $15$ and is a constituent of $S^{2}(W)\otimes W^{*}$ where $W$ is the 33-dimensional natural module. We can then calculate that $g$ has at least six distinct nontrivial eigenvalues on $V$.
@@
Denote by $E$ the number of distinct eigenvalues of $g$. Since $a\leq n/2$ and since nontrivial eigenspaces occur in orbits of size 33, it can be seen that $d+E$ is at most
@@
(attained when $g$ has trivial eigenspace of dimension $n/2$, three eigenspaces of dimension 11 and three eigenspaces of dimension (n/23)/3(n/2-3)/3), from which it also follows that $d\leq n^{2}/3-n+6$. If $G=\mathrm{SL}_{n}(q)$ then we deduce $|C_{G}(g)|^{3}<q^{n^{2}-3n+18}$, which is <|G:M|<|G:M| if $n\geq 15$. If $G=\mathrm{SU}_{n}(q^{1/2})$ then, setting $q_{0}=q^{1/2}$, by Lemma 7.11 we deduce $|C_{G}(g)|^{3}<q_{0}^{3(d+E)}\leq q_{0}^{n^{2}-3n+39}$, which is <|G:M|<|G:M| if $n\geq 24$.
Let us address the remaining cases (i.e. the case $n\leq 14$ for $G=\mathrm{SL}_{n}(q)$ and the case $n\leq 23$ for $G=\mathrm{SU}_{n}(q^{1/2})$). Assume first $V$ is $p$-restricted; the options are $n=6,10$ for $G=\mathrm{SL}_{n}(q)$; and $n=6,10,15,18,21$ for $G=\mathrm{SU}_{n}(q^{1/2})$. If $n=6$ then $V=S^{2}(W)$ (where $W$ is the natural 3-dimensional module), $g$ is regular and we have $|C_{G}(g)|<|G:M|^{1/3}$. If $n=10$ then $V=S^{3}(W)$; we see that $a=\dim(C_{V}(g))=1$ and so $|C_{G}(g)|<|G:M|^{1/3}$. If $n=15$ then it must be $a\leq 6$, which implies $d\leq 51$ and the result follows. If $n=18$ then $p=5$ and $V$ is a constituent of $S^{3}(W)\otimes W^{*}$. Since by the above the largest eigenspace on $S^{3}(W)$ has dimension at most 22, it follows that $a\leq 6$, which gives the conclusion. Finally, if $n=21$ then $a\leq 9$, so $d\leq 111=3\cdot 6^{2}+3$ and this is sufficient. Assume now $V$ is not $p$-restricted with $n\leq 23$; by [70], the possible embeddings are of type $\mathrm{SL}_{3}(q^{2})<\mathrm{SL}_{9}(q)$ and $\mathrm{SL}_{3}(q)<\mathrm{SU}_{9}(q^{1/2})$. In the first case, $a=3$ and every nontrivial eigenspace has dimension 11, so $d=15$ and the result follows. In the second case, there are three nontrivial eigenspaces of dimension 11, and three nontrivial eigenspaces of dimension 22, so again $d=15$ and $|C_{G}(g)|<|G:M|^{1/3}$ also in this case.
@@
Suppose $g$ is semisimple. Note that every prime divisor of $|g|$ is at least 77 and also we have $|N_{L}(\langle g\rangle)/C_{L}(g)|=3$. By [80, Theorem 1.2], we deduce that $g$ has at least 55 distinct eigenvalues. Since these are permuted in orbits of size 33, we deduce that $g$ has at least 66 distinct eigenvalues. In particular, by the same calculation as in the case $p\mid r$ we see that $d\leq n^{2}/3-n+6$ and $d+E\leq n^{2}/3-n+13$, where $E$ denotes the number of distinct eigenvalues of $g$. Denote also by $E_{2}$ the number of distinct irreducible factors of degree at least two of the characteristic polynomial of $g$. Next we upper bound $|C_{G}(g)|$ using Lemmas 7.10 and 7.11.
@@
Suppose finally $g$ is not semisimple; we apply Lemma 7.4. Letting $P$ be a Sylow $p$-subgroup of $\langle g\rangle$, we have that $V\downarrow P=V_{0}\oplus U$ where all Jordan blocks on $V_{0}$ have size $|P|\geq 7$, and $U$ is the sum of at most 33 indecomposable module of dimension $t$. If $t\geq 4$ then applying Lemma 7.8 with $C=0$ and $B=4$ we deduce $d\leq n^{2}/4$. In particular, the exact same computation as in Lemma 9.8 reduces to the case $n\leq 29$, for which we are done (as we are assuming $n\geq 41$).